OpenStudy (anonymous):

What is the equation of the line described below written in slope-intercept form? the line passing through point (-1, 5) and parallel to the line whose equation is x + y = 10 y = x - 4 y = x + 4 y = -x + 4 c

OpenStudy (anonymous):

y = -x + 10 The slope of that expression is -1. Hence, by point-slope form, noting that parallel lines must have the slopes of -1 [same slopes], and that the point for the new line is (-1,5)... y - 5 = -(x + 1) y - 5 = -x - 1 y = -x + 4

OpenStudy (anonymous):

so i was correct...XP

OpenStudy (anonymous):

Yes sir!

OpenStudy (anonymous):

can i have medal??